OpenStudy (anonymous):

1-sin(x/2)=0 on [0, 2pi]

OpenStudy (anonymous):

\[1-\sin(\frac{x}{2})=0 \rightarrow -\sin(\frac{x}{2})=0-1 \rightarrow \sin(\frac{x}{2})=1\] \[\sin(\frac{x}{2})=\sin(\frac{\pi}{2}) \rightarrow \frac{x}{2}=\frac{\pi}{2} \rightarrow x=\pi\]